# Asif Tauhid, Sidahmed Benabderrahmane, Mohamad Altrabulsi, Ahamed Foisal, Talal Rahwan

### v1

@@ II Related work
II-A APT Detection and System Provenance
@@
II-B Graph Neural Networks for Anomaly Detection
@@
II-C Association Rule and Pattern Mining for Security
@@ III Proposed Framework
In this section, we present our detection framework, which leverages a provenance database of process events and netflow activities.
III-A Overview
@@ III-D Graph Construction
III-E Graph Autoencoder Architecture
We employ a Graph Autoencoder (GAE) composed of a two-layer Graph Convolutional Network (GCN) encoder and an inner-product decoder. Using the node feature matrix $X\in\mathbb{R}^{n\times d}$ and the symmetrically normalized adjacency matrix $\hat{A}$, the encoder computes hidden representations as
$H^{(1)}=\mathrm{ReLU}\!\left(\hat{A}XW^{(0)}\right),$ (1)
$Z=\hat{A}H^{(1)}W^{(1)},$ (2)
where $W^{(0)}\in\mathbb{R}^{d\times 128}$ and $W^{(1)}\in\mathbb{R}^{128\times 64}$ are learnable weight matrices, and $Z\in\mathbb{R}^{n\times 64}$ contains the resulting node embeddings.
The decoder reconstructs the adjacency matrix via inner products between node embeddings:
$\hat{A}_{ij}=\sigma\!\left(z_{i}^{\top}z_{j}\right),$ (3)
where $\sigma(\cdot)$ is the sigmoid function and $z_{i}$ denotes the embedding of node $i$. The model is trained for unsupervised link prediction using a binary cross-entropy objective and optimized with Adam (learning rate $0.05$) for 200 epochs. At inference time, we score each node by its local reconstruction error, computed as the average discrepancy between observed and reconstructed edges in its neighborhood, and use this score for anomaly ranking.
III-F Rare Pattern Mining
@@
III-G Score Boosting Mechanism
@@
with $d(i)$ denoting the degree of node $i$ in the rare-pattern graph and $d_{\max}=\max_{j}d(j)$ the maximum degree. Intuitively, processes that share rare behavioral patterns with many other processes receive a larger bonus, reflecting stronger evidence of unusual co-occurrence structure.
We tune the boosting weight $\alpha$ over the set $\{0.1,0.5,1.0,2.0,5.0,10.0\}$ and select $\alpha=2.0$ based on validation performance.
III-H Workflow Summary
@@ IV-A Datasets
The datasets used in this study are derived from DARPA’s Transparent Computing (TC) program [^1] https://www.darpa.mil/program/transparent-computing and are processed using the ADAPT [^2] https://gitlab.com/adaptdata (Automatic Detection of Advanced Persistent Threats) ingester. The resulting corpus spans multiple operating systems, namely Android, Linux, BSD, and Windows, and includes two representative attack scenarios [7].
@@
The provenance records are first transformed into graph databases and then subjected to integration and deduplication. For learning and evaluation, each dataset is represented in a process-centric tabular form: each row corresponds to a process and is encoded as a high-dimensional Boolean feature vector, where a value of $1$ indicates the presence of a particular behavioral attribute or event associated with that process. Overall, the dataset collection comprises multiple operating systems and attack scenarios, yielding a diverse benchmark for provenance-based anomaly detection. Similar to prior work on DARPA TC, the resulting data are highly imbalanced, with attack-related processes forming a very small fraction of all observed processes.
@@
Figure 3 visualizes the process-level behavioral space for Linux data, after projecting all 272,376 processes (each represented by 299 features) onto the first two principal components. Although PC1 and PC2 together capture roughly 52% of the variance, the plot shows substantial overlap between normal and attack processes, with attack points dispersed across regions dominated by benign behavior rather than forming a clearly separable cluster. This highlights a key challenge of provenance-based detection under extreme class imbalance: malicious activity often manifests as subtle deviations that remain close to normal operating patterns in a low-dimensional view. The absence of a clean boundary in the PCA projection motivates our use of relational modeling and rare-pattern evidence—rather than relying on feature-space separation alone—to prioritize suspicious processes based on structural inconsistency and infrequent behavioral co-occurrences.
@@
Figure 3: PCA visualization of process behavioral feature space for Linux Dataset. We project all 272,376 processes, each described by 299 behavioral features, into two dimensions using Principal Component Analysis (PCA) for visualization. The first two components (PC1 and PC2) together capture approximately 52% of the total variance, highlighting the overall distribution of processes in the reduced space (with attack instances shown in red and normal instances in gray).
IV-B Evaluation Metric
Our goal is to rank processes such that truly malicious (or injected attack) processes appear as early as possible in the ordered list. We therefore adopt the normalized Discounted Cumulative Gain (nDCG), a standard metric for evaluating the quality of ranked retrieval results. Let $\pi$ denote the predicted ranking over $n$ processes, and let $\mathrm{rel}_{\pi(k)}\geq 0$ be the relevance of the item at rank $k$ (e.g., $\mathrm{rel}\in\{0,1\}$ for benign vs. attack). The Discounted Cumulative Gain at cutoff $K$ is defined as
$\mathrm{DCG@}K\;=\;\sum_{k=1}^{K}\frac{2^{\mathrm{rel}_{\pi(k)}}-1}{\log_{2}(k+1)}.$ (6)
To make results comparable across datasets and attack scenarios, DCG is normalized by the ideal DCG obtained from a perfect ranking:
$\mathrm{nDCG@}K\;=\;\frac{\mathrm{DCG@}K}{\mathrm{IDCG@}K},$ (7)
where $\mathrm{IDCG@}K$ is computed by sorting items in descending order of ground-truth relevance. nDCG takes values in $[0,1]$, with larger values indicating better ranking quality, and it explicitly emphasizes retrieving highly relevant instances at early ranks through logarithmic discounting.
IV-C Performance Results
@@
Table I reports the impact of incorporating rare-pattern evidence through the boosting weight $\alpha$. When $\alpha=0$, the model reduces to the baseline GAE that ranks processes purely using reconstruction-based deviations. As $\alpha$ increases from small values (e.g., $\alpha=0.1$) to moderate values (e.g., $\alpha\in\{0.5,1.0,2.0\}$), the ranking performance improves steadily, indicating that the rare-pattern signal provides complementary information beyond structural reconstruction alone. The best performance is obtained at $\alpha=2.0$, suggesting that a moderate emphasis on rare-pattern connectivity most effectively amplifies truly suspicious processes without overwhelming the baseline reconstruction signal. For larger values (e.g., $\alpha\geq 5.0$), performance saturates and slightly declines, consistent with the intuition that overly strong boosting may start to favor highly connected rare-pattern neighborhoods even when their structural deviations are weak. Overall, these results validate the benefit of integrating rare-pattern evidence: the hybrid score consistently improves upon the GAE-only detector and yields the most accurate prioritization when the two signals are well-weighted.
IV-D Comparison with Baseline Methods
Table II compares our boosted model against representative unsupervised baselines previously reported on the same DARPA TC E2 Linux/pandex setting (Scenario 1, ProcessAll context). The proposed GAE + Rare Pattern approach achieves the strongest anomaly-ranking performance among the listed single-method baselines. In particular, it improves over classical provenance anomaly detectors such as compression- and statistics-based approaches, and it also surpasses prior rule-mining baselines that rely on association-rule scoring. Notably, even without rare-pattern boosting, the baseline GAE remains competitive with the strongest rule-mining baseline and exceeds the remaining classical detectors, highlighting the effectiveness of graph reconstruction as a standalone unsupervised signal in provenance graphs.
@@ IV-D Comparison with Baseline Methods
Beyond aggregate performance, Table II also supports a practical deployment advantage: unlike approaches that rely on combining multiple independent detectors, our framework provides a single unified model in which rare-pattern evidence is incorporated directly into the scoring mechanism. This unification simplifies deployment while preserving interpretability, since high-scoring processes can be traced back to the rare co-occurrence patterns that contribute to their boosted score.
@@
TABLE I: Effect of rare-pattern boosting on nDCG.
@@
TABLE II: Comparison of nDCG scores on DARPA TC E2 Linux/pandex (Scenario 1, ProcessAll context).
Method Type nDCG
Our approach
GAE + Rare Pattern ($\alpha=2.0$) GNN + Mining 0.66
Baseline GAE GNN 0.54
Benabderrahmane et al.
VR-ARM Rule Mining 0.54
VF-ARM Rule Mining 0.13
Berrada et al.
OC3 Compression 0.41
AVF Statistical 0.29
@@ V Conclusion
This paper introduced a hybrid neuro-symbolic framework for provenance-based anomaly detection that combines graph representation learning with rare behavioral pattern mining. We construct a process similarity graph from provenance-derived features and learn normal relational structure using a Graph Autoencoder (GAE) trained via unsupervised link reconstruction. To improve sensitivity to infrequent but security-relevant behaviors, we mine rare co-occurrence patterns with Apriori and inject this evidence through a degree-based score boosting mechanism on a rare-pattern graph. The resulting approach leverages complementary notions of abnormality: structural deviations captured by graph reconstruction and symbolic rarity captured by mined patterns.
Experimental results on provenance traces from the DARPA Transparent Computing program demonstrate that rare-pattern guidance consistently strengthens reconstruction-based scoring and improves anomaly ranking in highly imbalanced settings. Beyond performance, the proposed framework retains interpretability by allowing high-risk processes to be associated with the rare patterns they support, providing actionable cues for investigation.
In future work, we plan to (i) integrate pattern evidence directly into the training objective through pattern-aware regularization or heterogeneous graph modeling, (ii) extend the framework to streaming or online settings with incremental pattern updates, and (iii) incorporate richer temporal and causal constraints to better capture multi-stage attack behaviors across diverse host environments.

### v2

@@ Abstract
Code: https://github.com/asiftauhid/gae_anomaly_detection
@@ II Related work
APT Detection and System Provenance:
@@
Graph Neural Networks for Anomaly Detection
@@
Association Rule and Pattern Mining for Security:
@@ III Proposed Framework
Overview:
@@
Graph Autoencoder Architecture:
We employ a Graph Autoencoder (GAE) composed of a two-layer Graph Convolutional Network (GCN) encoder and an inner-product decoder. Given the node feature matrix $X\in\mathbb{R}^{n\times d}$ and the symmetrically normalized adjacency matrix with self-loops $\hat{A}$, the encoder computes $H^{(1)}=\mathrm{ReLU}\!\left(\hat{A}XW^{(0)}\right)$ and $Z=\hat{A}H^{(1)}W^{(1)}$, where $W^{(0)}\in\mathbb{R}^{d\times h_{1}}$ and $W^{(1)}\in\mathbb{R}^{h_{1}\times h_{2}}$ are learnable parameters. The resulting embedding matrix $Z\in\mathbb{R}^{n\times h_{2}}$ provides a compact representation of process behavior that is used by the decoder to reconstruct graph connectivity.
The decoder reconstructs the adjacency matrix via inner products between node embeddings: $\hat{A}_{ij}=\sigma\!\left(z_{i}^{\top}z_{j}\right)$, where $\sigma(\cdot)$ is the sigmoid function and $z_{i}$ denotes the embedding of node $i$. The model is trained for unsupervised link prediction using a binary cross-entropy objective and optimized with Adam (learning rate $0.05$) for 200 epochs. At inference time, we score each node by its local reconstruction error, computed as the average discrepancy between observed and reconstructed edges in its neighborhood, and use this score for anomaly ranking.
Rare Pattern Mining:
@@
Score Boosting Mechanism:
@@
with $d(i)$ denoting the degree of node $i$ in the rare-pattern graph and $d_{\max}=\max_{j}d(j)$ the maximum degree. Intuitively, processes that share rare behavioral patterns with many other processes receive a larger bonus, reflecting stronger evidence of unusual co-occurrence structure. We tune the boosting weight $\alpha$ over the set $\{0.1,0.5,1.0,2.0,5.0,10.0\}$ and select $\alpha=2.0$ based on validation performance.
Workflow Summary:
@@ IV Experimental settings, results and analysis
The datasets used in this study are derived from DARPA’s Transparent Computing (TC) program [^1] https://www.darpa.mil/program/transparent-computing and are processed using the ADAPT [^2] https://gitlab.com/adaptdata (Automatic Detection of Advanced Persistent Threats) ingester. The resulting corpus spans multiple operating systems, namely Android, Linux, BSD, and Windows, and includes two representative attack scenarios [7]. The code and configuration files necessary to reproduce our experiments are publicly available in the project repository [^3] https://github.com/asiftauhid/gae_anomaly_detection.
@@
The provenance records are first transformed into graph databases and then subjected to integration and deduplication. For learning and evaluation, each dataset is represented in a process-centric tabular form: each row corresponds to a process and is encoded as a high-dimensional Boolean feature vector, where a value of $1$ indicates the presence of a particular behavioral attribute or event associated with that process. Similar to prior work on DARPA TC, the resulting data are highly imbalanced (BSD: 13 anomalies /75,903 benign; Linux: 8/17,569; Windows: 25/247,160; Android: 9/102), with attack-related processes forming a very small fraction of all observed processes.
@@
Figure 3 visualizes the process-level behavioral space for Linux data, after projecting all processes (each represented by 299 features) onto the first two principal components. Although PC1 and PC2 together capture roughly 52% of the variance, the plot shows substantial overlap between normal and attack processes, with attack points dispersed across regions dominated by benign behavior rather than forming a clearly separable cluster. This highlights a key challenge of provenance-based detection under extreme class imbalance: malicious activity often manifests as subtle deviations that remain close to normal operating patterns in a low-dimensional view. The absence of a clean boundary in the PCA projection motivates our use of relational modeling and rare-pattern evidence—rather than relying on feature-space separation alone—to prioritize suspicious processes based on structural inconsistency and infrequent behavioral co-occurrences.
@@
Figure 3: PCA visualization of process behavioral feature space for Linux Dataset. We project all processes, each described by 299 behavioral features, into two dimensions using Principal Component Analysis (PCA) for visualization. The first two components (PC1 and PC2) together capture approximately 52% of the total variance, highlighting the overall distribution of processes in the reduced space (with attack instances shown in red and normal instances in gray).
Evaluation Metric:
Although the area under the ROC curve (AUC) is widely used for binary classification, it can be less informative in provenance-based detection due to extreme class imbalance, where benign processes outnumber attack processes by several orders of magnitude. In such settings, a detector may achieve a strong AUC while still failing to prioritize the few truly malicious processes near the top of the output list, which is the operational requirement in security triage. We therefore evaluate performance from a ranking perspective, aiming to surface the most suspicious processes first. To measure ranking quality, we use the normalized Discounted Cumulative Gain (nDCG), a standard metric in information retrieval [5]. nDCG rewards methods that rank relevant (attack) processes early, while progressively discounting gains at lower ranks, reflecting the diminishing utility of late retrieval in practical investigations. The metric is normalized by the score of an ideal (perfect) ranking, yielding values between 0 and 1 and enabling fair comparison across datasets and attack scenarios.
Performance Results:
Figure 4: Training loss curve of RPG-AE (BSD dataset). The loss drops sharply during the first epochs, exhibits mild early oscillations, and then gradually stabilizes to a low plateau, indicating convergence.
@@
Figure 4 presents the training loss curve of RPG-AE (BSD data). The early fluctuations reflect initial optimization instability, while the later flat region suggests the model has learned a stable reconstruction of the normal graph structure. Table I (together with Figure 5) reports the impact of incorporating rare-pattern evidence through the boosting weight $\alpha$. When $\alpha=0$ (ablation), the model reduces to the baseline GAE that ranks processes purely using reconstruction-based deviations. As $\alpha$ increases from small values (e.g., $\alpha=0.1$) to moderate values (e.g., $\alpha\in\{0.5,1.0,2.0\}$), the ranking performance improves steadily, indicating that the rare-pattern signal provides complementary information beyond structural reconstruction alone. The best performance is obtained at $\alpha=2.0$, suggesting that a moderate emphasis on rare-pattern connectivity most effectively amplifies truly suspicious processes without overwhelming the baseline reconstruction signal. For larger values (e.g., $\alpha\geq 5.0$), performance saturates and slightly declines, consistent with the intuition that overly strong boosting may start to favor highly connected rare-pattern neighborhoods even when their structural deviations are weak. Overall, these results validate the benefit of integrating rare-pattern evidence: the hybrid score consistently improves upon the GAE-only detector and yields the most accurate prioritization when the two signals are well-weighted.
Comparison with Baseline Methods:
Table II summarizes the anomaly-ranking performance of our method against representative baselines on the same DARPA TC provenance benchmark. Our proposed RPG-AE consistently delivers strong performance across operating systems and achieves the best nDCG on Linux, Windows, and Android, while matching the top result on BSD. Specifically, rare-pattern guidance yields clear gains over the graph-only baseline (GNN), improving nDCG from 0.56 to 0.66 on Linux, from 0.48 to 0.62 on Windows, from 0.25 to 0.52 on BSD, and from 0.41 to 0.85 on Android. These improvements confirm that rare co-occurrence evidence complements reconstruction-based deviations and is particularly effective in highly imbalanced settings where attacks may be weakly separable in feature space.
Compared to prior families of detectors, RPG-AE also outperforms classical statistical/compression approaches and standard outlier detectors (e.g., Isolation Forest, OC-SVM, LOF), which generally lag behind on Linux/Windows/BSD and exhibit larger cross-OS variability. In addition, RPG-AE improves upon association-rule baselines (VR-ARM and VF-ARM), which can perform competitively in some operating systems but fail to generalize consistently across all environments (e.g., near-zero performance on Android). Overall, these results indicate that combining a unified graph autoencoding backbone with rare-pattern mining yields a robust and portable detector that generalizes across diverse operating systems.
@@
Beyond aggregate accuracy, Table II highlights a practical advantage of our design: unlike approaches that rely on multiple context-specific detectors or ensemble aggregation, RPG-AE provides a single end-to-end framework where symbolic rarity evidence is injected directly into the anomaly scoring function. This integration simplifies deployment and supports interpretability, since elevated scores can be traced back to the rare behavioral patterns that contributed to the final ranking.
@@
TABLE I: Effect of rare-pattern boosting on nDCG (with relative improvement).
@@
Figure 5: Effect of rare-pattern boosting on nDCG.
TABLE II: Comparison of nDCG scores on DARPA TC.
Method Type nDCG Score
Linux Windows BSD Android
RPG-AE (Ours) GNN+Rare patterns 0.66 0.62 0.52 0.85
GNN Graph Baseline 0.56 0.48 0.25 0.41
VR-ARM [5] Rule mining 0.54 0.61 0.36 0
VF-ARM [5] Rule mining 0.13 0.50 0.18 0
OC3 [7] Compression 0.41 0.43 0.38 0.82
AVF [7] Statistical 0.29 0.52 0.52 0.83
I-Forest[7] ML 0.17 0.12 0.13 0.32
OC-SVM[7] ML 0.17 0.14 0.13 0.32
LOF [7] ML 0.15 0.018 0.07 0.18
DeepOCCL [21] Deep Learning 0.20 0.26 0.22 0.75
@@ V Conclusion
This paper introduced a hybrid neuro-symbolic framework for provenance-based anomaly detection that combines graph representation learning with rare behavioral pattern mining. We construct a process similarity graph from provenance-derived features and learn normal relational structure using a GAE trained via unsupervised link reconstruction. To improve sensitivity to infrequent but security-relevant behaviors, we mine rare co-occurrence patterns with Apriori and inject this evidence through a degree-based score boosting mechanism on a rare-pattern graph. The resulting approach leverages complementary notions of abnormality: structural deviations captured by graph reconstruction and symbolic rarity captured by mined patterns. Experimental results on provenance traces from the DARPA TC program demonstrate that rare-pattern guidance consistently strengthens reconstruction-based scoring and improves anomaly ranking in highly imbalanced settings. Beyond performance, the proposed framework retains interpretability by allowing high-risk processes to be associated with the rare patterns they support, providing actionable cues for investigation. In future work, we plan to (i) integrate pattern evidence directly into the training objective through pattern-aware regularization or heterogeneous graph modeling, (ii) extend the framework to streaming or online settings with incremental pattern updates, and (iii) incorporate richer temporal and causal constraints to better capture multi-stage attack behaviors across diverse host environments.